OpenStudy (anonymous):

please help.... x^2+2x-5=0 solve by completing the square

OpenStudy (tkhunny):

Well, do it. x^2 + 2x + ______ = 5 + ______ Fill in the blanks.

OpenStudy (anonymous):

I dont know how?

OpenStudy (anonymous):

@tkhunny

OpenStudy (imstuck):

Separate the 5 from the rest of the eqatuion and put it on the other side of the equal sign. Take half the x term (half of 2 is 1) and square the 1 to get 1. Add the 1 to both sides of the equation, which gives you x^2 + 2x + 1 = 5 + 1. Now, what did you do to get the 1 in the constant position? Squared the 1. Right? So the left side of the equation looks like this: (x +1)^2 and the right side is a 6. (x + 1)^2 = 6. Now take the square root of both sides, leaving x + 1 = square root of 6. Subtract the 1 and that leaves x = -1 +/- sq rt 6. Do you get that?

OpenStudy (tkhunny):

You must exploit this principle. \((a+b)^{2} = a^{2} + 2ab + b^{2}\) That is all you are doing. Please note that you have shown no work on this thread. You MUST do better than that.